OpenStudy (anonymous):

Using g(x), Write a function that shows grams as a function of cups, when 0.5 cup of flour has a mass of 64 grams and 2 cups of flour has a mass of 256 grams

OpenStudy (anonymous):

@whpalmer4, do you understand this?

OpenStudy (whpalmer4):

Yes, I do. Anything else? :-)

OpenStudy (whpalmer4):

this reads like it is part of a bigger problem, however; is it possible to see the whole thing?

OpenStudy (anonymous):

2. While in the United States visiting her grandmother, Kendra copied the famous family apple pie recipe. When she returns to England, she plans on making the same apple pie for her friends. To do this, Kendra needs to convert the flour from cups to grams. She knows that 0.5 cup of flour has a mass of 64 grams and 2 cups of flour has a mass of 256 grams. A) Write a function that shows grams as a function of cups. Use g(x). @whpalmer4

OpenStudy (whpalmer4):

okay. I wanted to make sure that there wasn't some definition of g(x) that we would need to know about. There isn't. so, if 0.5 cups flour masses 64 grams, how much is the mass of 1.0 cups of flour?

OpenStudy (anonymous):

128

OpenStudy (whpalmer4):

Okay, good. Can you write a function of \(x\) called \(g(x)\) that converts \(x\) (in cups) to a mass in grams?

OpenStudy (whpalmer4):

apparently, you just have to multiply the number of cups by 128 grams/cup, wouldn't you agree?

OpenStudy (anonymous):

Well, my function was \[g(x) =0.5c\] but I dont think its right

OpenStudy (whpalmer4):

no, it isn't. what is \(c\)?

OpenStudy (anonymous):

C is the number of cups.

OpenStudy (whpalmer4):

and your function doesn't do anything with the argument \(x\), so it is going to return the same value every time

OpenStudy (whpalmer4):

the number of cups is the argument to the function, \(x\)

OpenStudy (whpalmer4):

you need in the function to do a calculation which will have as its result the number of grams of flour that is equivalent to \(x\) cups of flour

OpenStudy (anonymous):

So, maybe something like \[g(x) = c +something\]?

OpenStudy (anonymous):

Functions confuse me. Dx

OpenStudy (whpalmer4):

okay, 3 cups of flour. convert that to grams, and show your work.

OpenStudy (whpalmer4):

do you add, subtract, multiply, divide, take square roots or logarithms, what?

OpenStudy (anonymous):

I don't know..

OpenStudy (whpalmer4):

how many grams in 1 cup of flour?

OpenStudy (anonymous):

We said 128g

OpenStudy (whpalmer4):

okay, so 1 cup = how many grams?

OpenStudy (whpalmer4):

2 cups = how many grams? 3 cups = how many grams?

OpenStudy (anonymous):

1 cup = 128 2 cups = 192 3 cups = 256

OpenStudy (whpalmer4):

well, 1 of the 3 is correct. if 1 cup is 128 grams, shouldn't 2 cups be the same as 1 cup + 1 cup = 128 + 128 =

OpenStudy (anonymous):

I added 64 both times, so I gues I actually did 1 cups, 1.5 cups, and 2 cups.

OpenStudy (anonymous):

So, 1 cup = 128 2 cups = 256 3 cups = 384

OpenStudy (whpalmer4):

yes, indeed you did. so how much would 3 cups be, in grams?

OpenStudy (anonymous):

384g

OpenStudy (whpalmer4):

okay. how much would \(x\) cups be, in grams?

OpenStudy (anonymous):

g(x) = x + 128?

OpenStudy (whpalmer4):

how did you calculate how many grams 3 cups would be?

OpenStudy (anonymous):

I added the amount of 1 cup to itself twice.

OpenStudy (whpalmer4):

isn't it simply a straightforward multiplication of the number of cups * the number of grams per cup?!?

OpenStudy (whpalmer4):

\[x \text{ cups} * \frac{128\text{ grams}}{1\text{ cup}} = x \cancel{\text{ cups}}*\frac{128\text{ grams}}{\cancel{1\text{ cup}}} = 128x \text{ grams}\]

OpenStudy (anonymous):

Ohh! g(x) = c * 128?

OpenStudy (anonymous):

if C = number of cups

OpenStudy (whpalmer4):

there IS NO C!!!

OpenStudy (whpalmer4):

\(x\) is the number of cups. if \(c\) represented the number of cups, we would be asked to create \(g(c)\), not \(g(x)\)

OpenStudy (whpalmer4):

Yes, \(g(x) = 128x\) if \(x\) is the number of cups of flour